Car A and car B are travelling down a straight road.Car A is travelling at 15m/s and accelerating at 2m/s2.Car B starts from speed 0 but accelerates at 4m/s2.After how many seconds car A and B will be travelling at the same speed?


1
Expert's answer
2021-03-08T08:24:22-0500

Let's write the velocities of cars A and B:


"v_A=v_{0A}+a_{0A}t,""v_B=v_{0B}+a_{0B}t."

Cars A and B will be travelling at the same speed when "v_A=v_B":


"v_{0A}+a_{0A}t=v_{0B}+a_{0B}t,""15+2t=0+4t,""t=\dfrac{15\ \dfrac{m}{s}}{2\ \dfrac{m}{s^2}}=7.5\ s."